# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 8.1 Counterfactual Regret Minimization
The key concept in CFR is the counterfactual value, which measures how much better player $i$ would have done by taking action $a$ at information set $I$, assuming she played to reach $I$ but her opponent played according to $\sigma$? Formally:
@@
where $\pi^{\sigma}_{-i}(h)$ is the probability opponents reach $h$, and $\pi^{\sigma}(ha,z)$ is the probability of reaching terminal $z$ from $ha$. The counterfactual formulation weights by opponent reach $\pi^{\sigma}_{-i}(h)$ rather than joint reach $\pi^{\sigma}(h)$. This ensures non-zero updates even at rarely-visited information sets, avoiding the vanishing gradient problem that affects direct policy gradient methods in extensive-form games. Cumulative regret for action $a$ is $R^{T}(I,a)=\sum_{t=1}^{T}[v^{\sigma^{t}}(I,a)-v^{\sigma^{t}}(I)]$. CFR updates via regret matching:112112112Regret matching is an action selection rule where the probability of choosing action $a$ is proportional to the cumulative regret for not having chosen $a$ in the past (truncated at zero). Actions with high regret receive higher probability; actions with negative regret (having performed worse than average) receive zero probability.
@@
The average strategy $\bar{\sigma}^{T}$ converges to $\varepsilon$-Nash with $\varepsilon=O(|\mathcal{I}|\sqrt{|\mathcal{A}|}\Delta/\sqrt{T})$, where $\Delta$ is the range of payoffs (Zinkevich et al., 2008).
CFR+ (Tammelin, 2014) replaces standard regret matching with Regret Matching+, which truncates cumulative regrets to zero after each update rather than only at action selection. The update becomes $R^{T+1}(I,a)=\max(R^{T}(I,a)+r^{T+1}(I,a),\,0)$, where $r^{T+1}(I,a)=v^{\sigma^{T+1}}(I,a)-v^{\sigma^{T+1}}(I)$ is the instantaneous counterfactual regret. CFR+ also weights iteration $t$ by $t$ when computing the average strategy, giving later iterations, which reflect strategies closer to equilibrium, more influence. While vanilla CFR converges at $O(1/\sqrt{T})$, CFR+ empirically converges at $O(1/T)$, roughly squaring the number of useful digits of exploitability per unit of compute.113113113The $O(1/T)$ rate for CFR+ is empirically observed but not yet proven in full generality. Tammelin (2014) conjectured this rate based on extensive experiments across poker variants. CFR+ enabled Bowling et al. (2015) to essentially solve heads-up limit Texas hold’em, a game with $3.16\times 10^{17}$ states, the first non-trivial imperfect-information game played competitively by humans to be essentially solved. Their program Cepheus achieved exploitability below 1 mbb/g (milli-big-blind per game), small enough that a human lifetime of play would not suffice to distinguish it from a Nash equilibrium strategy.
@@ 8.2.1 Deep CFR
Weighting by iteration index $t^{\prime}$ gives more recent regret estimates higher importance, since later iterations reflect strategies closer to equilibrium. A separate network $\Pi_{\phi}$ approximates the average strategy, trained via weighted MSE on strategy samples with the same iteration weighting. The current strategy derives from regret matching, with $\sigma(I,a)\propto[V_{\theta}(I,a)]^{+}$.
@@ 8.2.2 Neural Fictitious Self-Play
NFSP (Heinrich and Silver, 2016) combines fictitious play114114114Fictitious play (Brown, 1951) is a classical learning rule where each player best-responds to the empirical distribution of opponents’ past actions. Under certain conditions (e.g., two-player zero-sum games), the time-average strategies converge to Nash equilibrium. (Brown, 1951) with deep Q-learning.115115115Deep Q-Network (DQN) (Mnih et al., 2015) approximates the Q-function with a neural network, using experience replay and a target network to stabilize training. The target network $\theta^{-}$ in the loss function is a delayed copy of the main network, updated periodically. Each player maintains two networks: a best-response network $Q_{\theta}$ trained via DQN, and an average strategy network $\bar{\pi}_{\phi}$ trained via supervised learning on the best-response actions:
@@ 8.2.3 Poker Results
These methods achieved superhuman performance in poker. Deep CFR attained exploitability of 37 mbb/g (milli-big-blinds per game)116116116Exploitability measures how far a strategy is from Nash equilibrium, defined as the maximum expected gain an adversary could achieve by best-responding. Zero exploitability means the strategy is unexploitable (Nash). In poker, exploitability is measured in milli-big-blinds per game (mbb/g). in heads-up flop hold’em (Brown et al., 2019). Libratus defeated top human professionals in heads-up no-limit hold’em using nested subgame solving with CFR (Brown and Sandholm, 2018). Pluribus extended to six-player no-limit hold’em (Brown and Sandholm, 2019). The game tree of heads-up no-limit hold’em contains $\sim 10^{161}$ states; these results demonstrate that CFR-based methods scale to economically relevant game sizes.
@@ 8.3 The Coase Conjecture
The durable goods monopoly problem has two distinct cases that yield qualitatively different equilibria. In the “gap case,” the seller’s cost is strictly below all buyer valuations, so trade is guaranteed. In the “no-gap case,” cost equals or exceeds the lowest buyer valuation, so some buyers may never trade. The gap case admits a unique stationary equilibrium; the no-gap case admits multiple equilibria with more complex dynamics. I study the gap case, which has a clean theoretical benchmark.
Coase (1972) conjectured that a durable goods monopolist117117117A durable good provides utility over multiple periods (e.g., a car, appliance, or software license) rather than being consumed immediately. Unlike non-durable goods, durable goods create intertemporal competition, as the seller at time $t$ competes with her own future self at $t+1$. loses market power when buyers are patient. Unable to commit to future prices, the seller competes with her future self, eroding rents. Stokey (1981) showed that a monopolist facing rational consumers cannot price discriminate intertemporally; Bulow (1982) proved that in the no-gap case, the monopolist prefers renting to selling. Gul et al. (1986) provided the definitive formalization. With a continuum of buyer types and continuous time, as the inter-offer interval $\Delta t\to 0$, price converges to marginal cost instantaneously and all trade occurs at $t=0$.
@@ 8.3.1 Model
A seller with zero cost faces a buyer with private valuation $v\in\{v_{L},v_{H}\}$, where $\Pr(v=v_{H})=\pi$. In each period $t=1,2,\ldots$, the seller posts price $p_{t}$; the buyer accepts or rejects. Upon acceptance at $t$, the seller receives $\delta^{t-1}p_{t}$ and the buyer receives $\delta^{t-1}(v-p_{t})$, where $\delta\in(0,1)$ is the common discount factor. The game ends upon acceptance or after $T$ periods.
This is the gap case, where seller cost (0) is strictly below the lowest buyer valuation ($v_{L}$), guaranteeing trade in equilibrium.
8.3.2 Theoretical Benchmark
The screening price $P^{*}(\delta)$ makes the high-type buyer indifferent between accepting now and waiting:
@@ 8.3.2 Theoretical Benchmark
For $\pi<\pi^{*}$, the seller pools (offer $v_{L}$ immediately). For $\pi>\pi^{*}$, the seller screens (offer $P^{*}$, then $v_{L}$ if rejected).118118118In a screening equilibrium, the seller uses price to separate buyer types. High-value buyers accept immediately at a high price, while low-value buyers reject and receive a lower offer later. In a pooling equilibrium, the seller offers a single price that all types accept, sacrificing price discrimination for immediate sale.
The Coase conjecture manifests as $\delta\to 1$, where the screening price $P^{*}(\delta)\to v_{L}$, so the seller cannot extract surplus from high types.
@@ 8.3.3 Computational Results
I model the bargaining game as an extensive-form game and apply CFR. Parameters: $v_{L}=100$, $v_{H}=200$, $T=2$ periods. The seller’s information set is the rejection history; the buyer’s information set includes her private type.
@@
Notes: P(Screen) is the probability the seller offers $P^{*}=150$ in period 1. Theory column gives the predicted probability (0 for pooling, 1 for screening). NashConv is the sum of exploitabilities. 5,000 CFR iterations per configuration.
Table 12 reports results from varying $\pi$ at fixed $\delta=0.5$. The algorithm recovers the phase transition at $\pi^{*}=0.5$. For $\pi<0.5$, the seller pools with probability approaching 1; for $\pi>0.5$, the seller screens with probability approaching 1. The transition is sharp.
A $\delta$-sweep at $\pi=0.7$ confirms the screening price formula: $P^{*}(\delta)=200-100\delta$ with zero error across $\delta\in[0.1,0.9]$. As $\delta$ increases, screening becomes unprofitable (Coase conjecture); at $\delta=0.75$, the seller switches to pooling.
8.3.4 Validation
Four stress tests verify the results.119119119(1) Awkward primes $(v_{L},v_{H})=(37,83)$: converged to $P^{*}=55$ (theory: 55.4). (2) Information leak: single seller information set at root. (3) Grid shift with 150 removed: used 145 (99.4%). (4) 3-period game: found $P_{1}=120$ vs. theory 136. The 3-period result merits explanation. At $P_{1}=136$, the high buyer is exactly indifferent: surplus $v_{H}-136=64=\delta(v_{H}-120)$. If the buyer randomizes, the seller prefers $P_{1}=120$ (revenue 108 vs. 86.4 if buyer rejects 136). CFR found the equilibrium robust to tie-breaking.
@@ 8.4 Discussion
CFR computed equilibria matching the theoretical predictions of Gul et al. (1986) without any knowledge of the economic structure. The algorithm discovered that (1) the seller screens when $\pi$ is high, pools when $\pi$ is low, with threshold at $\pi^{*}=0.5$; (2) the screening price satisfies the high-buyer indifference condition; (3) as $\delta\to 1$, screening collapses (Coase conjecture). These properties emerged from regret minimization alone.
9 Real-Time Reinforcement Learning and Economic Modelling
In this chapter we focus on “online” or “real-time” reinforcement learning where and agent learns “live” or “in production” and not in a “offline” or simulation environment. This setting makes “exploration” a real cost, and the agent is forced to balance exploration vs exploitation. This setting is particularly relevant for economic applications, where the agent is typically a firm learning about demand and the cost of exploration is lost revenue. I focus on the problem of dynamic pricing, where a seller learns to set prices for products with unknown demand. This problem has been studied extensively in the economics and machine learning literatures, and it provides a rich testbed for understanding how structural assumptions about demand affect algorithmic learning.
Rothschild (1974) posed pricing under demand uncertainty as a two-armed bandit problem. His insight was that a myopic seller can get stuck at a suboptimal price forever, because exploiting the currently best-looking price generates no information about alternatives. This chapter focuses on papers that develop these types of explore-exploit models and tries to quantify the value of structural modeling in online dynamic pricing.
A seller faces $T$ customers in sequence, setting a price for each and observing only whether the customer bought. The seller does not observe the customer’s willingness to pay. Regret measures the total revenue gap between the seller’s pricing policy and the best fixed price in hindsight. If the optimal fixed price earns revenue $r^{*}$ per customer, a policy with regret $R(T)$ earns approximately $Tr^{*}-R(T)$ in total. The central question in dynamic pricing is how fast $R(T)$ can shrink, relative to $T$, as the seller accumulates purchase data.
This chapter surveys the theoretical and empirical literature on bandit-based dynamic pricing. I begin with foundational results that establish the baseline regret rates under progressively stronger assumptions on demand. I then examine more recent research, each illustrating a distinct economic insight. A comparison table collects these results and I conclude with a simulation study replicating the Misra et al. (2019) pricing experiment.
@@ 9.1.1 No Structure on Demand
Kleinberg and Leighton (2003) study the simplest version of the problem. Customers arrive with valuations drawn i.i.d. from an unknown distribution on $[0,1]$, and the seller posts a price from a continuous set. The demand curve $D(p)=\Pr(v\geq p)$ is unknown; the only feedback is whether each customer bought. The seller’s goal is to minimize regret against the single price $p^{*}$ that maximizes $p\cdot D(p)$. Kleinberg and Leighton prove that the minimax regret is $\Theta(\sqrt{T})$.120120120The upper bound, $O(\sqrt{T\log T})$, discretizes $[0,1]$ into $K=\lceil(T/\log T)^{1/4}\rceil$ prices and runs the UCB1 algorithm of Auer et al. (2002). The lower bound, $\Omega(\sqrt{T})$, constructs a family of demand curves parameterized by the location of the optimal price $p^{*}\in[0.3,0.4]$. The key tension: posting prices far from $p^{*}$ is informative about demand but costly in revenue; posting prices near $p^{*}$ is cheap but uninformative. Resolving this tension costs at least $\Omega(\sqrt{T})$ in cumulative revenue. In concrete terms, after 10,000 customers the seller loses roughly 100 customers’ worth of revenue to the uncertainty in demand. No algorithm can do better without imposing structure on $D$. For adversarial valuations (chosen by a worst-case opponent rather than drawn from a fixed distribution), the minimax regret rises to $\Theta(T^{2/3})$, achieved by the Exp3 algorithm on a discretized price grid. I focus on the stochastic setting throughout this chapter.
@@ 9.1.2 Parametric Demand
Broder and Rusmevichientong (2012) consider a parametric demand model $d(p;z)=\Pr(V\geq p)$, where $z\in\mathbb{R}^{n}$ is an unknown parameter governing the demand curve. The seller observes binary purchase decisions and updates a maximum likelihood estimate of $z$. Under standard regularity conditions (bounded demand, unique optimal price, smooth revenue function), the minimax regret is again $\Theta(\sqrt{T})$.121121121The lower bound (Theorem 3.1 of Broder and Rusmevichientong (2012)) constructs a linear demand family where all demand curves pass through the same point at the optimal price $p^{*}(z_{0})$. Observing purchases at this price provides no information about $z$. An MLE-Cycle policy that interleaves dedicated exploration rounds with greedy pricing achieves the matching upper bound $O(\sqrt{T})$ (Theorem 3.6). Parametric structure alone does not break the $\sqrt{T}$ barrier.
The picture changes under a “well-separated” condition requiring that every price in the feasible set is informative about the demand parameter. Formally, the Fisher information $I(p,z)$ is bounded below by $c_{f}>0$ for all prices $p$ and parameters zz.122122122This means no two demand curves $d(p;z)$ and $d(p;z^{\prime})$ cross on the pricing interval, so every purchase observation distinguishes the two hypotheses. In the lower-bound family of Theorem 3.1, the curves all cross at $p=1$, which is why the $\sqrt{T}$ floor emerges. Under well-separation, a pure greedy policy (MLE-Greedy, which simply prices at $p^{*}(\hat{z}_{t})$ in each period) achieves $O(\log T)$ regret (Theorem 4.8). After 10,000 customers, the revenue loss is only about 9 customers’ worth. When every price teaches the seller something, dedicated exploration rounds become unnecessary.

### 15 Mar 2026 v2

@@ 10.2 Counterfactual Regret Minimization
Extensive-form games require a different approach. CFR bypasses equilibrium selection: instead of computing Nash equilibria at each step, it minimizes cumulative regret and lets the time-averaged strategy converge to equilibrium.
@@
The counterfactual value of action $a$ at information set $I$ is
@@
where $\pi^{\sigma}_{-i}(h)$ is the probability opponents reach $h$, and $\pi^{\sigma}(ha,z)$ is the probability of reaching terminal $z$ from $ha$. The counterfactual formulation weights by opponent reach $\pi^{\sigma}_{-i}(h)$ rather than joint reach $\pi^{\sigma}(h)$, ensuring non-zero updates even at rarely-visited information sets. Cumulative regret for action $a$ is $R^{T}(I,a)=\sum_{t=1}^{T}[v^{\sigma^{t}}(I,a)-v^{\sigma^{t}}(I)]$. CFR updates via regret matching:158158158Regret matching is an action selection rule where the probability of choosing action $a$ is proportional to the cumulative regret for not having chosen $a$ in the past (truncated at zero). Actions with high regret receive higher probability; actions with negative regret (having performed worse than average) receive zero probability.
@@
The average strategy $\bar{\sigma}^{T}$ converges to $\varepsilon$-Nash with $\varepsilon=O(|\mathcal{I}|\sqrt{|\mathcal{A}|}\Delta/\sqrt{T})$, where $\Delta$ is the range of payoffs (Zinkevich et al., 2008).159159159An $\varepsilon$-Nash equilibrium is a strategy profile where no player can improve her expected payoff by more than $\varepsilon$ through unilateral deviation. As $\varepsilon\to 0$, this converges to an exact Nash equilibrium.
CFR+ (Tammelin, 2014) replaces standard regret matching with Regret Matching+, which truncates cumulative regrets to zero after each update rather than only at action selection. The update becomes $R^{T+1}(I,a)=\max(R^{T}(I,a)+r^{T+1}(I,a),\,0)$, where $r^{T+1}(I,a)=v^{\sigma^{T+1}}(I,a)-v^{\sigma^{T+1}}(I)$ is the instantaneous counterfactual regret. CFR+ also weights iteration $t$ by $t$ when computing the average strategy. While vanilla CFR converges at $O(1/\sqrt{T})$, CFR+ empirically converges at O(1/T)O(1/T).160160160The $O(1/T)$ rate for CFR+ is empirically observed but not yet proven in full generality. Tammelin (2014) conjectured this rate based on extensive experiments across poker variants. CFR+ enabled Bowling et al. (2015) to essentially solve heads-up limit Texas hold’em, a game with $3.16\times 10^{17}$ states, the first non-trivial imperfect-information game played competitively by humans to be essentially solved. Their program Cepheus achieved exploitability below 1 mbb/g (milli-big-blind per game).
@@ 10.3.1 Deep CFR
Weighting by iteration index $t^{\prime}$ gives more recent regret estimates higher importance. A separate network $\Pi_{\phi}$ approximates the average strategy, trained via weighted MSE on strategy samples with the same iteration weighting. The current strategy derives from regret matching, with $\sigma(I,a)\propto[V_{\theta}(I,a)]^{+}$.
@@ 10.3.2 Neural Fictitious Self-Play
NFSP (Heinrich and Silver, 2016) combines fictitious play161161161Fictitious play (Brown, 1951) is a classical learning rule where each player best-responds to the empirical distribution of opponents’ past actions. Under certain conditions (e.g., two-player zero-sum games), the time-average strategies converge to Nash equilibrium. (Brown, 1951) with deep Q-learning.162162162Deep Q-Network (DQN) (Mnih et al., 2015) approximates the Q-function with a neural network, using experience replay and a target network to stabilize training. The target network $\theta^{-}$ in the loss function is a delayed copy of the main network, updated periodically. Each player maintains two networks: a best-response network $Q_{\theta}$ trained via DQN, and an average strategy network $\bar{\pi}_{\phi}$ trained via supervised learning on the best-response actions:
@@ 10.3.3 Poker Results
These methods achieved superhuman performance in poker. Deep CFR attained exploitability of 37 mbb/g163163163Exploitability measures how far a strategy is from Nash equilibrium, defined as the maximum expected gain an adversary could achieve by best-responding. Zero exploitability means the strategy is unexploitable (Nash). In poker, exploitability is measured in milli-big-blinds per game (mbb/g). in heads-up flop hold’em (Brown et al., 2019). Libratus defeated top human professionals in heads-up no-limit hold’em using nested subgame solving with CFR (Brown and Sandholm, 2018). Pluribus extended to six-player no-limit hold’em (Brown and Sandholm, 2019). The game tree of heads-up no-limit hold’em contains $\sim 10^{161}$ states; these results demonstrate that CFR-based methods scale to economically relevant game sizes.
@@ 10.4 The Coase Conjecture
The durable goods monopoly is a canonical extensive-form bargaining problem with private information: the seller does not know the buyer’s valuation.
Coase (1972) conjectured that a durable goods monopolist164164164A durable good provides utility over multiple periods (e.g., a car, appliance, or software license) rather than being consumed immediately. Unlike non-durable goods, durable goods create intertemporal competition, as the seller at time $t$ competes with her own future self at $t+1$. loses market power when buyers are patient. Unable to commit to future prices, the seller competes with her future self, eroding rents. As the inter-offer interval shrinks ($\delta\to 1$), price collapses to marginal cost and all surplus is extracted by buyers. Gul et al. (1986) formalized this for the gap case, where the seller’s cost is strictly below all buyer valuations, guaranteeing trade and a unique stationary equilibrium.165165165Stokey (1981) showed that a monopolist facing rational consumers cannot price discriminate intertemporally; Bulow (1982) proved that in the no-gap case, the monopolist prefers renting to selling. The no-gap case admits multiple equilibria with more complex dynamics.
@@ 10.4.1 Model
A seller with zero cost faces a buyer with private valuation $v\in\{v_{L},v_{H}\}$, where $\Pr(v=v_{H})=\pi$. In each period $t=1,2,\ldots$, the seller posts price $p_{t}$; the buyer accepts or rejects. Upon acceptance at $t$, the seller receives $\delta^{t-1}p_{t}$ and the buyer receives $\delta^{t-1}(v-p_{t})$, where $\delta\in(0,1)$ is the common discount factor. The game ends upon acceptance or after $T$ periods. Parameters: $v_{L}=100$, $v_{H}=200$, $T=2$ periods.
10.4.2 Equilibrium Analysis
The screening price $P^{*}(\delta)$ makes the high-type buyer indifferent between accepting now and waiting.
@@ 10.4.2 Equilibrium Analysis
For $\pi<\pi^{*}$, the seller pools (offers $v_{L}$ immediately). For $\pi>\pi^{*}$, the seller screens (offers $P^{*}$, then $v_{L}$ if rejected).166166166In a screening equilibrium, the seller uses price to separate buyer types: high-value buyers accept immediately at a high price, while low-value buyers reject and receive a lower offer. In a pooling equilibrium, the seller offers a single price that all types accept. The gap case ($0=c<v_{L}$) guarantees trade in equilibrium. The Coase conjecture manifests as $\delta\to 1$, where $P^{*}(\delta)\to v_{L}$ and the seller cannot extract surplus from high types.
@@ 10.4.3 Computational Results
I model the bargaining game as an extensive-form game and apply CFR.167167167The game tree has two information sets for the seller (indexed by rejection history) and two for the buyer (indexed by private type). CFR runs for 5,000 iterations per parameter configuration; NashConv (sum of exploitabilities) measures convergence.
@@
Notes: P(Screen) is the probability the seller offers $P^{*}=150$ in period 1. Theory column gives the predicted probability (0 for pooling, 1 for screening).
Table 20 reports results from varying $\pi$ at fixed $\delta=0.5$. CFR recovers the sharp phase transition at $\pi^{*}=0.5$: pooling below, screening above. A $\delta$-sweep at $\pi=0.7$ confirms the screening price formula $P^{*}(\delta)=200-100\delta$ with zero error across $\delta\in[0.1,0.9]$; at $\delta\approx 0.75$, the seller switches to pooling as patient buyers erode the screening premium, consistent with the Coase conjecture.168168168Four stress tests were conducted: (1) awkward primes $(v_{L},v_{H})=(37,83)$, converging to $P^{*}=55$ (theory: 55.4); (2) information leak with a single seller information set at the root; (3) grid shift with 150 removed, recovering 145 (99.4%); (4) 3-period game, finding $P_{1}=120$ versus theoretical 136. The 3-period result reflects a tie-breaking equilibrium: at $P_{1}=136$ the high buyer is exactly indifferent, so the seller prefers $P_{1}=120$ (revenue 108 versus 86.4 if rejected).
@@ 10.5 Discussion
Stochastic-game Q-learning and CFR target complementary game classes: simultaneous-move games with observable payoffs and extensive-form games with private information, respectively. The simulations confirm convergence to known equilibria in both settings without encoding economic structure into the algorithms.
11 Bandits and Dynamic Pricing
This chapter focuses on online reinforcement learning (Section 2), where exploration is a real cost and the agent must balance exploration against exploitation. A seller faces $T$ customers in sequence, setting a price for each and observing only whether the customer bought.169169169Rothschild (1974) posed pricing under demand uncertainty as a two-armed bandit problem. His insight was that a myopic seller can get stuck at a suboptimal price forever, because exploiting the currently best-looking price generates no information about alternatives. The seller does not observe the customer’s willingness to pay. Regret, the total revenue gap between the seller’s policy and the best fixed price in hindsight, is the central measure: if the optimal fixed price earns $r^{*}$ per customer, a policy with regret $R(T)$ earns $Tr^{*}-R(T)$ in total. The central question is how fast $R(T)$ shrinks as the seller accumulates purchase data, and how structural assumptions about demand affect this rate.
@@ 11.1.1 No Structure on Demand
Kleinberg and Leighton (2003) study the simplest version of the problem. Customers arrive with valuations drawn i.i.d. from an unknown distribution on $[0,1]$, and the seller posts a price from a continuous set. The demand curve $D(p)=\Pr(v\geq p)$ is unknown; the only feedback is whether each customer bought. The seller’s goal is to minimize regret against the single price $p^{*}$ that maximizes $p\cdot D(p)$. Kleinberg and Leighton prove that the minimax regret is $\Theta(\sqrt{T})$.170170170The upper bound, $O(\sqrt{T\log T})$, discretizes $[0,1]$ into $K=\lceil(T/\log T)^{1/4}\rceil$ prices and runs the UCB1 algorithm of Auer et al. (2002a). The lower bound, $\Omega(\sqrt{T})$, constructs a family of demand curves parameterized by the location of the optimal price $p^{*}\in[0.3,0.4]$. The key tension: posting prices far from $p^{*}$ is informative about demand but costly in revenue; posting prices near $p^{*}$ is cheap but uninformative. Resolving this tension costs at least $\Omega(\sqrt{T})$ in cumulative revenue. UCB1 (Auer et al., 2002a) selects the price maximizing $\hat{\mu}_{p_{k}}(t)+\sqrt{2\ln t/N_{p_{k}}(t)}$, where $\hat{\mu}_{p_{k}}(t)$ is the empirical mean profit and $N_{p_{k}}(t)$ the number of trials; the second term is an exploration bonus that shrinks as a price is tried more, implementing the principle of optimism in the face of uncertainty. In concrete terms, after 10,000 customers the seller loses roughly 100 customers’ worth of revenue to the uncertainty in demand. No algorithm can do better without imposing structure on $D$. For adversarial valuations (chosen by a worst-case opponent rather than drawn from a fixed distribution), the minimax regret rises to $\Theta(T^{2/3})$, achieved by the Exp3 algorithm171171171Exp3 (Auer et al., 2002b) maintains a weight $w_{k}(t)$ for each price, selecting $p_{k}$ with probability proportional to $w_{k}(t)$ and updating the chosen price’s weight by $\exp(\eta\hat{r}_{k,t})$ where $\hat{r}_{k,t}$ is the revenue importance-weighted by the selection probability; because no model of demand is assumed, the guarantee holds against an adversary who chooses valuations after observing the algorithm. on a discretized price grid. I focus on the stochastic setting throughout this chapter.
@@ 11.1.2 Parametric Demand
Broder and Rusmevichientong (2012) consider a parametric demand model $d(p;z)=\Pr(V\geq p)$, where $z\in\mathbb{R}^{n}$ is an unknown parameter governing the demand curve. The seller observes binary purchase decisions and updates a maximum likelihood estimate of $z$. Under standard regularity conditions (bounded demand, unique optimal price, smooth revenue function), the minimax regret is again $\Theta(\sqrt{T})$.172172172The lower bound (Theorem 3.1 of Broder and Rusmevichientong (2012)) constructs a linear demand family where all demand curves pass through the same point at the optimal price $p^{*}(z_{0})$. Observing purchases at this price provides no information about $z$. An MLE-Cycle policy that interleaves dedicated exploration rounds with greedy pricing achieves the matching upper bound $O(\sqrt{T})$ (Theorem 3.6). Parametric structure alone does not break the $\sqrt{T}$ barrier.
The picture changes under a “well-separated” condition requiring that every price in the feasible set is informative about the demand parameter. Formally, the Fisher information $I(p,z)$ is bounded below by $c_{f}>0$ for all prices $p$ and parameters zz.173173173This means no two demand curves $d(p;z)$ and $d(p;z^{\prime})$ cross on the pricing interval, so every purchase observation distinguishes the two hypotheses. In the lower-bound family of Theorem 3.1, the curves all cross at $p=1$, which is why the $\sqrt{T}$ floor emerges. Under well-separation, a pure greedy policy (MLE-Greedy, pricing at $p^{*}(\hat{z}_{t})$ each period) achieves $O(\log T)$ regret (Theorem 4.8), because every price is informative and dedicated exploration rounds become unnecessary.
Figure 18: Revenue curves $r(p)=r^{*}-k(p-p^{*})^{2}$ for four demand curvatures $k\in\{0.5,1.0,2.0,3.5\}$. All models agree at the optimal price $p^{*}=5$. Within the shaded exploration zone, the curves are nearly indistinguishable, so playing prices near $p^{*}$ is uninformative about the demand parameter.